OpenStudy (eyust707):

what is the potential function of -yi + xj what is the potential function of -yi + xj @Mathematics

OpenStudy (anonymous):

What math class is this for?

OpenStudy (eyust707):

calc 3

OpenStudy (eyust707):

we need some function whos partil with respect to x is -y and whose partial with respect to y is x

OpenStudy (eyust707):

whats the startegy to find such a function

OpenStudy (anonymous):

Fun class... LOL umm ok have you done multiple integration or iterated integrals?

OpenStudy (eyust707):

yes

OpenStudy (eyust707):

but i am pretty sure we do not use iterated integrals to find a vector fields potential

OpenStudy (eyust707):

We need a function that satisfies: \[\frac{\partial f}{\partial x} = -y\] and \[\frac{\partial f}{\partial y} = x\]

OpenStudy (anonymous):

hmm... i dont remember exactly, I know you have to integrate both the i and j components, then solve for the "constant of integration" and then recombine them to obtain a final answer...

OpenStudy (eyust707):

yea.. sounds correct heres an example of one..

OpenStudy (eyust707):

i think the potential would be f(x,y) = -yx + x^2